# Sample Observed Effects: Enumeration, Randomization and Generalization

### v5

@@ 3 Sampling Treatment-Background Assignments
We will start with a definition for sampling sequences, and subsets of permutations $\Pi_{n}(Z/X)$, when we can make repeated measurements for the observed variables, $X$, in the DGP. Eq.(3) formulated a definition of effects for a single fixed factor $a$, where backgrounds were $\Pi_{n}(Z/\{a\})$. It is generalized here for samples $X$. These definitions rely on the following assumptions, which are somewhat typical,
@@
We also make two simplifying assumptions,
@@
    Large-scale data. The sample $X$ includes both causes and spurious factors. In our applications of interest, such as large-scale economic, genetic and medical data, samples typically contain both.
  2. 2.
    Non-strongly antagonistic unobserved interactions. When samples contain true causes Principle (1) holds (Appendix.B). Many times, however, causes are unobserved but we still want to find in-sample surrogates that closely reproduce their root cause effects. In this case, Principle (1) might not hold when the effects of unobserved causes (i.e., observed in the same sample gap) strongly interact and can cancel each other.
Sampling sequences can be interpreted mathematically in several ways. We choose to think of them as permutations $\pi:b(t)$ of the set $K$ of integer sample gap indices. Imagine our effect observations to be discrete and equally spaced points in time. If an effect, sampled at time $t$, is next sampled at time $t^{\prime}$, we write $b(t)=t^{\prime}$. If $t$ is an integer then let $t\bmod q$ denote the remainder when $t$ is divided by $q$. Finally, we let $[t]_{q}$ be a shorthand for the set of possible remainders $\{0,1,2,...,q-1\}$ in this division. We call each factor $k$ in $[t]_{q}$ an individual sample gap index, $u_{k}$, Eq.(2). Remember that permutations can have fixation points (positions whose order do not change). Each individual sample effect observation is thus associated with a subset of unobserved factors $u_{k}\subseteq U$, Eq.(2), and we say, in this case, that the effect is observed in the background of $u_{k}$. Considering multiple sample factors, each permutation defines, in turn, distinct treatment-background assignments. Regularities across assignments in $U{-}X$ systems define their treatment-background mechanism.
@@
Fig.2(a) illustrate this concept in a sample $X_{3}=\{a,b,c\}$ with a fixed order statistic, $a<b<c$, and changing $U_{4}=\{u_{1},u_{2},u_{3},u_{4}\}$ (slow sampling). The diagram shows the transition of backgrounds under repeated observations of effects with a fixed-order $a<b<c$. Backgrounds are partitioned in this example as $\{\{u_{1}\},\{u_{2},u_{3}\},u_{4}\}$. Each arrow indicates the sequential mapping of unobserved factors (arrow labels) across sample gaps (intervals). See  [Buhler and Graham., c2004.] for analogous combinatorial concepts.
Effect backgrounds $Z/X$ are not necessarily constant due to distinct frequencies, correlations and periodicities among $U$ and $X$ factors. A key definition in this article is
@@
In Fig.2(a), the top and middle diagrams are CF non-increasing. The bottom is CF non-increasing for a period of $2$. In Example (1), the sequences $\mathbf{ab}u_{1}u_{2}u_{3}\,,\allowbreak u_{1}\mathbf{ab}u_{2}u_{3}\,,u_{1}u_{2}\mathbf{ab}u_{3},\,\allowbreak u_{1}u_{2}u_{3}\mathbf{ab}.$ are CF non-increasing, but so are any sequences with cyclic permutations of in-sample factors (i.e., subsequent rows of sample squares) under the condition that the unobserved background is fixed.
@@ 3.1 Background Enumeration
When $q\gg m$, the constant $c_{1}$ describes the rate of increase in the counts of out-sample factors, and $c_{2}$ of in-sample factors. We first consider fast sampling, the case where $\Delta_{(k)}m>\Delta_{(k)}q$ in specific sampling periods. In this case, in-sample effects are observed for all factors in constant backgrounds. The constant $c_{2}$ describe ’phase’ differences that, when $c_{2}$ and $q$ are co-prime, generate samples with the cyclic permutations of $X$ (individual square rows). Systems under fast sampling are EV-increasing for a constant sampling rate when $c_{1}/q$ is integral,
@@
Consider, for example, $c_{1}/q=2$. For a fixed gap index $k$, it takes $2^{t}\times k(\bmod 2n-1)=k$ time for $u_{k}$ to get back to its original position  [Diaconis et al., 1983], and for all background factors to be enumerated111as result of the recursion $(2(2...(2k(\bmod 2n-1))...)))$ for each gap position $k$.. This can be restated as
@@
for a initial position $k=1$. This condition is known to only ’circulate’ the set of factors, $b(t)$, and lead to no randomization  [Diaconis et al., 1983, Diaconis and Fulman, 2023]. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}<u_{4}\;\}$, for example, where $q=4$ and $c_{1}=2$, linear extensions of type $\mathbf{a}u_{1}u_{2}\mathbf{b}u_{3}u_{4}$ can generate EV-increasing sequences with a constant sampling rate, unlike $\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b}u_{4}$.
@@ 3.2 Background Randomization
We next consider the case of in-sync and slow sampling, $q\gg m$. Both in-sync and slow sampling can be seen as allowing the multiplicative and integral constant $c_{1}$ to take random values. In the in-sync case, $c_{1}$ takes approximately common $\Delta_{(k)}q$ values across all $k$, $0<k\leq m$. It then describes background-treatment assignments where the $q$ confounders are equally distributed across sample gaps, $\Delta_{(k)}q\times m\approx q$. In Example (1), we exemplified how sample effect estimation depends on the extensions between posets $G=\{G(X),G(U)\}$. We then described them as a partition of $U$ induced by sample gaps in Eq.(2). More specifically, we can describe this relationship with a multinomial sample gap distribution, $p=\{p_{1},p_{2},...,p_{m}\}$.
Let $p=\{p_{1},p_{2},...,p_{m}\}$ then be a sample gap background assignment. Consider a set of subsequent unobserved $U=Z\backslash X$ backgrounds, Definition (1). Generalizability requires effects to be valid across any $\Pi_{n}(Z/X)$, irrespective of $p$. Imagine we keep $p$ constant from time $t=0$ to $\delta t$ while allowing $\Pi_{n}(Z/X)$ to change. Can we say anything about whether the backgrounds $\pi_{0}$ and $\pi_{\delta t}$ are random-like (i.e., the realization of $U$ in the first is as ’separated’ to the second as independent samples of a uniform distribution), $\pi_{0}\perp\pi_{t}$? We can describe this relationship more precisely by fixing an initial permutation $\pi_{0}$ and counting the number of inversions, or pairs of elements that are in the reverse order, from $\pi_{0}$ at the subsequent time $t$. Fix $p$ and two backgrounds $\pi_{0},\pi_{1}\in\Pi_{n}(Z/X)$ separated by only one inversion, for example. The likelihood of reaching $\pi_{1}$ from $\pi_{0}$ after one time step is high in this case. With increasing inversions in $\pi_{0}$, there should be some period $\delta t$ where that likelihood becomes comparable to that of a uniform distribution, where the expected number of inversions generated at each time is $\binom{m}{2}$. The expected number of permutation inversions, $I(\pi_{0},t)$, required for that condition (’randomization’) with a fixed distribution $p$, initial background $\pi_{0}$, and under i.i.d. sampling  [Fulman, 1998] is
@@
which underlies the known cut-off limit for markov chains  [Diaconis, 1996, Bayer and Diaconis, 1992]. As consequence, under random backgrounds (and where $U{-}X$ sampling is i.i.d.), effect generalizability requires $2^{t}$ effect samples. This is a stochastic generalization of Definition (5) and the previous case of $c_{1}=2$ on fast $U-X$ systems, $\nicefrac{{\Delta_{(k)}}}{{q}}\sim Bin(q,0.5)^{-1}$. Another trivial example has probability mass $p=\{1,0,0,...\}$ concentrated in a single gap, which, as expected, generates no inversions. This leads to the limit $|\Pi_{n}|=0$ which describes samples with no EV increasing sequences. A sample with $p$ close to $p=\{1/q,1/q,1/q,...\}$ ($q\geq m$) is capable of asymptotically enumerating $|\Pi_{n}|=q!$ orders of unobserved causes in the least time $t$, Eq.(2). This condition is associated with $U{-}X$ independence, and, departures from this condition with more restrictive requirements on sampling times or sample sizes. The number of EV-increasing permutations, for a given $t$, is decreasing as we depart the last limit (random $U$) and approach the previous (deterministic $p=\{1,0,0,0,...\}$). These cases are then associated with sample sizes of, respectively, order $\mathcal{O}(log_{2}t)$ and infinite for generalizable effect observations. For intermediary and more typical scenarios, what matters is the particular distribution $p$ of confounders across a sample’s gaps, Eq.(11). Notice that $\binom{m}{2}$ in Eq.(11) is the number of effects observation in a single square.
@@
Fig.2(b) illustrates this limit for the simulations in Sect.7 Experiments, which include samples with unobserved, correlated and non-stationary factors. The variation separation distance (y-axis) is a measure of randomness formulated as a distance to an uniform distribution  [Bayer and Diaconis, 1992, Diaconis, 1996]. The panel shows the distance between a fixed background $\pi_{0}$ and the subsequent in samples where $p$ is stationary (according to a Kendall-tau test, permutations and inversions are exhaustively enumerated and counted using the tests and procedures described next, distances where normalized to their per-factor range for display). The upper panel shows the same for other samples where $p$ is not stationary (according to Kendall-tau). These combinatorial considerations suggest limits where some guarantees on effect background randomization can be made. In systems with $q\gg m$, they guarantee that effects are observed in new backgrounds after each formulated period and, thus, that effects are observed under EV-increasing sampling sequences.
The proposed condition is only valid for a stationary $p$ and initial $\pi_{0}$. We associate EV-CF sequences, instead, with $m$ sampling sequences with the same properties but observed under $m$ in-sample cyclic permutations (i.e., with period $m\times\delta t$). The consequence of the later is that effects for all factors are observed across the same backgrounds for all in-sample factors $X$. These i.i.d. and permutation inversion conditions are further discussed in Sect.4 Square Enumeration and Gap Randomization Tests. In the context of effect estimation, Eq.(11) relates slow sampling to in-sync sampling (as visualized by squares) and the conditions under which they can lead to samples with generalizable effects. They implement the stipulated gap distributions in Eq.(8) and lead to the proposed solutions here for causal effect estimation with increasing generalizability,
@@ 4 Square Enumeration and Gap Randomization Tests
Both the randomization conditions in the last section and Principle (1) are related to permutation inversions. Inversions appear in common tests for i.i.d. sampling (’stationarity’) in ordinal data, such as in Kendall-tau and Man-Whitney tests. Any pair of observations $(x_{i},y_{i})$ and $(x_{j},y_{j})$, where $i<j$, are said to be ’concordant’ if the sort order of $(x_{i},x_{j})$ and $(y_{i},y_{j})$ agrees: that is, if either both $x_{i}>x_{j}$ and $y_{i}>y_{j}$ holds or both $x_{i}<x_{j}$ and $y_{i}<y_{j}$; and be ’discordant’ otherwise. The number of discordant pairs between $x$ and $y$ is equal to the number of inversions that permutes the $y$-sequence into the same order of $x$. The statistic can also be seen as a generalised measure of a classifier’s separation power for more than two classes  [Hand and Till, 2001].
@@
[…]
@@ 5 Related Work
Non-parametric tests (like permutation tests) can often only be used with strict assumptions on the completeness of samples. The previous discussions suggested the need to permute (unobserved) effect backgrounds when estimating effects. According to our view, due to the frequent partial observability of samples, the challenge for estimation then becomes how to establish in-sample conditions indicating when backgrounds (1) have not changed, or, (2) that they have changed enough such that any association between treatment and backgrounds of two effect observations or units nearly vanished. We suggested that a way to establish such conditions, when the number of unobserved factor is large, is to seek assurances of randomness from established theory, such as from Ergodic and Combinatorial Randomization theory  [Diaconis, 1996, Diaconis and Fulman, 2023].
@@
Concerns about robustness and predictive performance can, in part, be seen as the motivation for more recent views of causality based on model-performance invariance  [Peters et al., 2016, Magliacane et al., 2017, Buehlmann, 2020]. This is typically the invariance of a black-box model’s performance - and not of an effect, or, difference in outcomes, in a population, as taken here. The unit of analysis here, $\Delta y$, is much simpler (being an observation and not an estimate). Furthermore, on the present view, it is not the minimum variance of effects that matters, but the minimum variance of effects in face of the maximum variance of extraneous factors. The approach here addresses the latter conditions. Finally, we believe the widespread success of the counterfactual definition is largely due to its simplicity, and simplicity has been a central guideline in the present work.
@@
The issue of the Extenal Validity (EV) of effects has been addressed in  [Pearl and Bareinboim, 2014, 2011, Correa and Bareinboim, 2020]. Authors assume a known ’selection’ variable set $S$ that ’may represent all factors by which populations may differ or that may threaten the transport of conclusions between populations’ when training parametric models on non-experimental data. They then, assuming causal graphs for all other variables are known or inferred for both population, propose graphical conditions for EV based on Pearl’s graphical calculus. Assuming all this prior knowledge is known, the proposed conditions may give assurances on the transportability of ’observational findings’ (parameter estimation). Often the previous differences are not known. We take a more practical starting point, where every variable is potentially a relevant difference, no independence relations are assumed, and all results follow from in-sample effect differences and probabilities. We assume that subsets of ’population’ factors are observed ($X$) but another set of factors ($U$) is completely unobserved, as typical in supervised training tasks. We provide a quantifiable measure of EV (not a set of ideal graphical and independence conditions) for samples. In our view, the EV of effects is an uncertainty measure, Eq.(4), and not an isolated issue. The solution is formulated at the level of individual effect observations, non-parametrically, and in the counterfactual framework. At the same time, it is difficult to think of issues of external validity separate of issues of selection bias (which is the central issue in scientific research and Experimental Design)  [Cox, 2009, Montgomery, 2001] and randomization. We consider whether connections to Combinatorial Randomization might be helpful. The alternative combinatorial model proposed allow us to discuss critical issues - difficult to address, even if only theoretically, in the previous - like in-sample and out accuracy trade-offs, selection bias, prediction-interpretability tradeoffs, empirical results, and sample size requirements.

### v6

@@ 3 Sampling Treatment-Background Assignments
We will start with a definition for sampling sequences, assuming we can make repeated measurements for the observed variables in the DGP, $X$, but have no control over variables $U$. Eq.(3) formulated a definition of effects for a single factor $a$, where backgrounds were $\Pi_{n}(Z/\{a\})$. It is generalized here for samples $X$, $\Pi_{n}(Z/X)$. These definitions rely on the following assumptions, which are somewhat typical,
@@
We also make the simplifying assumption,
@@
    Non-strongly antagonistic unobserved interactions. When samples contain true causes Principle (1) holds (Appendix.B). Many times, however, causes are unobserved but we still want to find in-sample surrogates that closely reproduce their root cause’s effects. In this case, Principle (1) might not hold when the effects of unobserved causes (i.e., observed in the same sample gap) strongly interact and can cancel each other.
Sampling sequences can be interpreted mathematically in several ways. We choose to think of them as permutations $\pi:b(t)$ of the set $K$ of integer sample gap indices. Imagine our effect sample observations to be discrete and equally spaced points in time. If an effect, sampled at time $t$, is next sampled at time $t^{\prime}$, we write $b(t)=t^{\prime}$. If $t$ is an integer then let $t\bmod q$ denote the remainder when $t$ is divided by $q$. Finally, we let $[t]_{q}$ be a shorthand for the set of possible remainders $\{0,1,2,...,q-1\}$ in this division. We call each number $k$ in $[t]_{q}$ the sample gap assignment for an unobserved factors $u\subseteq U$, Eq.(2). Remember that permutations can have fixation points (positions whose order do not change). Each individual sample effect observation is consequently associated with a subset of unobserved factors $u\subseteq U$, Eq.(2), and we say, in this case, that the effect is observed in the background of $u$. Considering multiple sample factors, each permutation defines, in turn, distinct treatment-background assignments. Regularities across assignments define treatment-background mechanisms in $U{-}X$ systems.
@@
Fig.2(a) illustrates this concept in a sample $X_{3}=\{a,b,c\}$ with a fixed order statistic, $a<b<c$, and changing $U_{4}=\{u_{1},u_{2},u_{3},u_{4}\}$ (slow sampling). The diagram shows the transition of backgrounds under repeated observations of effects with the fixed sample order $a<b<c$. Backgrounds are partitioned as $\{\{u_{1}\},\{u_{2},u_{3}\},u_{4}\}$ in this example. Each arrow indicates the sequential mapping of unobserved factors (arrow labels) across sample gaps (intervals). See  [Buhler and Graham., c2004.] for analogous combinatorial concepts.
Effect backgrounds $Z/X$ are not necessarily constant across repeated effect observations due to distinct frequencies, correlations and periodicities among $U$ and $X$ factors. A key definition in this article is
@@
In Fig.2(a), the top and middle diagrams are CF non-increasing. The bottom is CF non-increasing for a period of $2$. In Example (1), the sequences $\mathbf{ab}u_{1}u_{2}u_{3}\,,\allowbreak u_{1}\mathbf{ab}u_{2}u_{3}\,,u_{1}u_{2}\mathbf{ab}u_{3},\,\allowbreak u_{1}u_{2}u_{3}\mathbf{ab}.$ are CF non-increasing, but so are any sequences with cyclic permutations of in-sample factors (i.e., subsequent rows of sample squares) under the condition that the unobserved background is fixed during the sample period.
@@ 3.1 Background Enumeration
We first consider fast sampling, the case where $\Delta_{(1)}m>\Delta_{(1)}q$ in specific sampling periods. In this case, in-sample effects are observed for all factors in constant backgrounds. The constant $c_{2}$ describe ’phase’ differences that, when $c_{2}$ and $q$ are co-prime, generate samples with the cyclic permutations of $X$ (individual square rows). Systems under fast sampling are EV-increasing for a constant sampling rate when $c_{1}/q$ is integral,
@@
Consider, for example, $c_{1}/q=2$. For a fixed gap assignment $k$, it takes $2^{t}\times k(\bmod 2n-1)=k$ time for $u_{k}$ to get back to its original position  [Diaconis et al., 1983], and for all background factors to be enumerated111as result of the recursion $(2(2...(2k(\bmod 2n-1))...)))$ for each gap position $k$.. This can be restated as
@@
for a initial position $k=1$. This condition is known to only ’circulate’ the set of treatment-background assignments, $b(t)$, and lead to no randomization  [Diaconis et al., 1983, Diaconis and Fulman, 2023]. For poset chains $\{G(X),G(U)\}=\{\;\mathbf{a}<\mathbf{b},\;\;u_{1}<u_{2}<u_{3}<u_{4}\;\}$, for example, where $q=4$ and $c_{1}=2$, linear extensions of type $\mathbf{a}u_{1}u_{2}\mathbf{b}u_{3}u_{4}$ can generate EV-increasing sequences with a constant sampling rate, unlike $\mathbf{a}u_{1}u_{2}u_{3}\mathbf{b}u_{4}$.
@@ 3.2 Background Randomization
We next consider the case of in-sync and slow sampling, $\Delta_{(1)}m\leq\Delta_{(1)}q$. Both in-sync and slow sampling can be seen as allowing the multiplicative and integral constant $c_{1}$ to take random values. In the in-sync case, $c_{1}$ takes approximately common $\Delta_{(k)}q$ values across all $k$, $0<k\leq m$. It then describes background-treatment assignments where the $q$ confounders are equally distributed across sample gaps, $\Delta_{(k)}q\times m\approx q$, across time. In Example (1), we exemplified how sample effect estimation depends on the extensions between posets $G=\{G(X),G(U)\}$. We then described them as partitions of $U$ induced by sample gaps, Eq.(2). We now describe this relationship with a multinomial sample gap distribution, $p=\{p_{1},p_{2},...,p_{m}\}$, counting the number of distinct confounders assigned to a gap $k$, $\Delta_{(k)}q$, after multiple sampling cycles.
Imagine we keep $p$ constant from time $t=0$ to $\delta t$ while allowing $\Pi_{n}(Z/X)$ to change. Can we say anything about whether the backgrounds $\pi_{0}$ and $\pi_{\delta t}$ are random-like (i.e., the realization of $U$ in the first is as ’separated’ to the second as independent samples of a uniform distribution), $\pi_{0}\perp\pi_{t}$? We can describe this relationship more precisely by fixing an initial permutation $\pi_{0}$ and counting the number of inversions, or pairs of elements that are in the reverse order, from $\pi_{0}$ at the subsequent time $t$. Fix $p$ and two backgrounds $\pi_{0},\pi_{1}\in\Pi_{n}(Z/X)$ separated by only one inversion, for example. The likelihood of reaching $\pi_{1}$ from $\pi_{0}$ after one time step is high in this case. With increasing inversions in $\pi_{0}$, there should be some period $\delta t$ where that likelihood becomes comparable to that of a uniform distribution, where the expected number of inversions generated at each time is $\binom{m}{2}$. The expected number of permutation inversions, $I(\pi_{0},t)$, required for that condition (’randomization’) with a fixed distribution $p$, initial background $\pi_{0}$, and under i.i.d. sampling  [Fulman, 1998] is
@@
which underlies the known cut-off limit for markov chains  [Diaconis, 1996, Bayer and Diaconis, 1992]. As consequence, for $m=2$ and random treatment-background assignments ($p_{1}{=}p_{2}{=}0.5$ and $U{-}X$ sampling i.i.d.), effect generalizability requires in the order of $\mathcal{O}(log_{2}t)$ effect samples. This is a stochastic generalization of Definition (5) and the previous case of $c_{1}=2$ on fast $U-X$ systems, $\nicefrac{{\Delta_{(k)}}}{{q}}\sim Bin(q,0.5)^{-1}$. Another trivial example has probability mass $p=\{1,0,0,...\}$ concentrated in a single gap, which, as expected, generates no inversions. This leads to the limit $|\Pi_{n}|=0$ which describes samples with no EV increasing sequences. A sample with $p$ close to $p=\{1/q,1/q,1/q,...\}$ ($q\geq m$) is capable of asymptotically enumerating $|\Pi_{n}|=q!$ orders of unobserved causes in the least time $t$, Eq.(2). This condition is associated with $U{-}X$ independence, and, departures from this condition with more restrictive requirements on sampling times or sample sizes. The number of EV-increasing permutations, for a given $t$, is decreasing as we depart the last limit (random $U$) and approach the previous (deterministic $p=\{1,0,0,0,...\}$). For intermediary and more typical scenarios, what matters is the particular distribution $p$ of confounders across a sample’s gaps, Eq.(11). Notice that $\binom{m}{2}$ in Eq.(11) is the number of effects observation in a single square.
@@
Fig.2(b) illustrates this limit for the simulations in Sect.7 Experiments, which include samples with unobserved, correlated and non-stationary factors. The variation separation distance (y-axis) is a measure of randomness formulated as a distance to an uniform distribution  [Bayer and Diaconis, 1992, Diaconis, 1996]. The panel shows the distance between a fixed background $\pi_{0}$ and the subsequent in samples where $p$ is stationary (according to a Kendall-tau test, permutations and inversions are exhaustively enumerated and counted using the tests and procedures described next, distances were normalized to their per-factor range for display). The upper panel shows the same for other samples where $p$ is not stationary (according to Kendall-tau). These combinatorial considerations suggest limits where some guarantees on effect background randomization can be made. In systems with $q\gg m$, they indicate that effects will be observed in new backgrounds after each formulated period and, thus, that effects are observed under EV-increasing sampling sequences.
The proposed condition is only valid for a stationary $p$ and initial $\pi_{0}$. We associate EV-CF sequences, instead, with $m$ sampling sequences with the same properties but observed under $m$ in-sample cyclic permutations (i.e., with period $m\times\delta t$). The consequence of the later is that effects for all factors are observed across the same backgrounds for all in-sample factors $X$. These i.i.d. and permutation inversion conditions are further discussed in Sect.4 Square Enumeration and Gap Randomization Tests. In the context of effect estimation, Eq.(11) relates slow sampling to ideal in-sync sampling (as visualized by squares) and the conditions under which they can lead to samples with generalizable effects. They implement the stipulated gap distributions in Eq.(8) and lead to the proposed solutions here for causal effect estimation with increasing generalizability,
@@ 4 Square Enumeration and Gap Randomization Tests
Both the randomization conditions in the last section and Principle (1) are related to permutation inversions. Inversions appear in standard tests for i.i.d. sampling (’stationarity’) in ordinal data, such as in Kendall-tau and Man-Whitney tests. Any pair of observations $(x_{i},y_{i})$ and $(x_{j},y_{j})$, where $i<j$, are said to be ’concordant’ if the sort order of $(x_{i},x_{j})$ and $(y_{i},y_{j})$ agrees: that is, if either both $x_{i}>x_{j}$ and $y_{i}>y_{j}$ holds or both $x_{i}<x_{j}$ and $y_{i}<y_{j}$; and be ’discordant’ otherwise. The number of discordant pairs between $x$ and $y$ is equal to the number of inversions that permutes the $y$-sequence into the same order of $x$. The statistic can also be seen as a generalised measure of a classifier’s separation power for more than two classes  [Hand and Till, 2001].
@@
$\begin{cases}\;\mathbb{I}_{C}\Big{[}\,\Delta^{2}y_{(k)}(X{=}x)\,\Big{]}_{t}=0,\\ \;\mathbb{I}_{C}\Big{[}\,\Delta y_{(k)}(X{=}x)\,\Big{]}_{t}>2\times\Big{(}p_{1}^{2}+p_{2}^{2}+...+p_{m}^{2}\Big{)}^{-t}\end{cases},$ (13)
@@ 5 Related Work
Non-parametric tests (like permutation tests) can often only be used with strict assumptions on sample completeness. The previous discussions suggested the need to permute (unobserved) effect backgrounds when estimating effects. According to our view, due to the frequent partial observability of samples, the challenge for estimation then becomes how to establish in-sample conditions indicating when backgrounds (1) have not changed, or, (2) that they have changed enough such that any association between treatment and backgrounds of two effect observations or units have nearly vanished. We suggested that a way to establish such conditions, when the number of unobserved factor is large, is to seek assurances of randomness from established theory, such as from Combinatorial Randomization theory  [Diaconis, 1996, Diaconis and Fulman, 2023].
@@
Concerns about robustness and predictive performance can, in part, be seen as the motivation for more recent views of causality based on model-performance invariance  [Peters et al., 2016, Magliacane et al., 2017, Buehlmann, 2020]. This is typically the invariance of a black-box model’s performance - and not of an effect, or, difference in outcomes, in a population, as taken here. The unit of analysis here, $\Delta y$, is much simpler (being an observation and not an estimate). Furthermore, on the present view, it is not the minimum variance of effects that matters, but the minimum variance of effects in face of the maximum variance of extraneous factors. The approach here formulates and addresses the latter conditions. Finally, we believe the widespread success of the counterfactual definition is largely due to its simplicity, and simplicity has been a central guideline in the present work.
@@
The issue of the Extenal Validity (EV) of effects has been addressed parametrically in  [Pearl and Bareinboim, 2014, 2011, Correa and Bareinboim, 2020]. Authors assume a known ’selection’ variable set $S$ that ’may represent all factors by which populations may differ or that may threaten the transport of conclusions between populations’ when training parametric models on non-experimental data. They then, assuming causal graphs for all other variables are known or inferred for both populations, propose graphical conditions for EV based on Pearl’s graphical calculus. Assuming all this prior knowledge is known, the proposed conditions may give assurances on the transportability of ’observational findings’ (parameter estimation). Often the previous differences are not known. We take a more practical starting point, where every variable is potentially a relevant difference, no independence relations are assumed, and all results follow from in-sample effect differences and probabilities. We assume that subsets of ’population’ factors are observed ($X$) but another set of factors ($U$) is completely unobserved, as typical in supervised training tasks. We provide a quantifiable measure of EV (not a set of ideal graphical and independence conditions) for samples. In our view, the EV of effects is an uncertainty measure, Eq.(4), and not an isolated issue. The solution is formulated at the level of individual effect observations, non-parametrically, and in the counterfactual framework. At the same time, it is difficult to think of issues of external validity separate of issues of selection bias (which is the central issue in scientific research and Experimental Design)  [Cox, 2009, Montgomery, 2001] and randomization. We consider whether connections to Combinatorial Randomization might be helpful. The alternative combinatorial model proposed allow us to discuss critical issues - difficult to address, even if only theoretically, in the previous - like in-sample and out accuracy trade-offs, selection bias, prediction-interpretability tradeoffs, empirical results, and sample size requirements.